OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line through point (-3, 4) and perpendicular to a line that has slope 2/5

OpenStudy (amorfide):

Okay, if you have any equation of a line

OpenStudy (amorfide):

okay I cant draw

OpenStudy (amorfide):

Any equation of a line y=2x/5 +3 the gradient (slope) is 2/5 if you have a line perpendicular to this line, the slope is -1/m in this case, m is 2/5 solve to work out your new gradient the equation of a line is y-y1=m(x-x1) where y1 is the y coordinate, x1 is the x coordinate

OpenStudy (anonymous):

i know but i dont get how to solve it ???

OpenStudy (amorfide):

okay so your perpendicular slope is -1/(2/5) which gives -5/2 so m=-5/2 y-y1=m(x-x1) substitute the coordinate values into y1and x1, where y1 is y coordinate, x1 is x coordinate so y-4=-5/2(x+3) simplify, get into form y=mx+c

OpenStudy (ranga):

They want the answer in standard form. So rearrange it to look like: Ax + By = C

OpenStudy (anonymous):

so 4x+-5=3

OpenStudy (anonymous):

is that right?

OpenStudy (amorfide):

y-4=-5/2(x+3) multiply by 2 2y-8=-5(x+3) expand the bracket 2y-8=-5x-15 2y+5x=-7

OpenStudy (anonymous):

Oh! lol i had the eqaution all backwards oops , but hey thanks @amorfide

OpenStudy (anonymous):

:)